# Generalized compression and compressive search of large datasets

### v1

@@ Abstract
1 Introduction
@@
In [13], we introduced a novel hierarchical clustering algorithm, and in [4], we introduced novel algorithms for efficient, exact $k$-NN search on large datasets. In this paper, we extend these previous works by introducing a novel hierarchical compression algorithm, panCAKES (pan- CLAM-Accelerated K-NN Entropy Scaling Search), which uses the indexing approach from [13] and [4] to enable efficient, exact $k$-NN and $\rho$-NN search on compressed data. We stress that panCAKES’ compression is aligned with, rather than orthogonal to, speedups provided by our indexing approach. Our hierarchical cluster tree allows for both efficient compression and search on the compressed dataset.
@@
2 Methods
2.1 Building the Tree
The compression approach presented in this paper relies on the divisive hierarchical clustering algorithm introduced in [4] and described in Algorithm 1.
@@ 2.1 Building the Tree
Note that Algorithm 1 does not necessarily produce a balanced tree. In fact, if we assume the manifold hypothesis–the notion that high-dimensional data collected from constrained generating phenomena typically only occupy a low-dimensional manifold within their embedding space–we expect anything but a balanced tree. The varying sampling density in different regions of the manifold and the low dimensional “shape” of the manifold itself will cause the tree to be unbalanced. The only case in which we would expect a balanced tree is if the dataset were uniformly distributed, e.g. in a $d$-dimensional hyper-sphere. We discuss additional ways that dataset structure informs properties of the cluster tree in Section 4.1.
2.2 Compression
@@
Of course, for massive datasets, radii at shallow depths of the cluster tree are not small, and as discussed in Section 4.1, they are not guaranteed to decrease after the first partition. Thus, unitary compression may not be the most efficient compression approach for all clusters.
@@ 3.5 MovieLens-10M
4 Results
4.1 Scaling Behavior of Cluster Radii
@@
To do this, we must first introduce the concept of fractal dimensionality of a dataset. We assume the manifold hypothesis (refer to [24] or Section 2.1); in other words, we assume that our dataset is embedded in a $D$-dimensional space, but that the data only occupy a $d$-dimensional manifold, where $d\ll D$. We refer to $d$ as the fractal dimensionality of the dataset.
@@
In this case, Algorithm 1 (Partition) will select a maximally distant pair of points to use as poles; without loss of generality, say we select two points at the extrema one of the $d$ axes. After one partition, the maximum pairwise distance along that axis will have been halved, i.e. bounded above by $R$. The next recursive partition will select another of the $d$ axes, and once again, the distance along that axis will be bounded above by $R$
Thus, after at most $d$ partitions, the maximum pairwise distance along each axis will be bounded above by $R$. The overall (i.e., not restricted to one axis) maximum pairwise distance will be bounded above by $R\sqrt{2}$ by for two instances that lie at the extrema of different axes. See Figure 3 for an example.
Figure 3: Scaling behavior of radii on a two dimensional; disk of uniformly distributed points representing the worst-case scenario for a two-dimensional distribution. The root cluster $C_{0}$ has a center $o_{0}$ and radius radius $R_{0}$. After one application of Partition (Algorithm 1), we have a child cluster $C_{1}$, with radius $R_{1}$ and center $o_{1}$. The right triangle formed by $o_{0}$, $o_{1}$, and $y_{+}$ in $C_{1}$ shows that $R_{0}<R_{1}$. Hence, the radius of a child cluster can be larger than that of its parent. However, after another application of Algorithm 1, we have consumed both orthogonal axes, as shown in $C_{2}$. Now, clearly $R_{2}<R_{0}$.
@@
4.2 Compression Upper Bound
In this section, we provide an upper bound on the cost of compression in order to illustrate how the cost changes with respect to LFD and depth in the tree. As in Section 4.1, we assume the worst-case distribution of data and that the local fractal dimension is uniform.
To illustrate the relationship between the LFD and the cost of compression, we introduce the notion of a “stride” in the tree. If $L$ is the LFD of the dataset, we can think of a stride as an $\lceil L\rceil$-depth moving window in the tree. For example, the first stride includes the all the clusters at depth 1 (root only) through depth $\lceil L\rceil$. The second stride includes the clusters at depths $\lceil L\rceil$ through $2\lceil L\rceil$. Hereafter, the ceiling is implied when LFD is non-integral; we use simply $L$ to mean $\lceil L\rceil$.
@@ 4.2 Compression Upper Bound
To compute $T_{R}$, we compute the recursive cost within each stride as a function of the stride index $i$, and then sum up over the number of strides $S$. Each edge (to a child cluster from its parent) in stride $i$ represents one encoding between a pair of cluster centers. By our analysis in Section 4.1, and by our choice of a distance function under which distance values are proportional to the memory cost of the encoding, we can use $r_{i}$, the radius of the shallowest cluster in stride $i$, as an upper bound for the memory cost of each such encoding.
@@
The first factor in Equation 7 uses the radii scaling behavior guarantee from Section 4.1 and the second factor is the number of non-center points in leaves at depth $S\cdot L$ (recall that we are assuming a balanced clustering).
@@
Figure 4: Compression cost breakdown assuming LFD = 3 and stride = 1. The leaf clusters within the stride (i.e., clusters at depth 4), shaded red, are compressed unitarily, Their ancestors, connected by green edges, are compressed using recursively. The recursive cost $T_{R}$ is the sum of the costs of encoding the green edges. The unitary cost $T_{U}$ is the sum of the costs of encoding all non-center points in the red clusters.
4.3 Compression Ratios
Table 1 reports the compression ratios of panCAKES and gzip. We observe that panCAKES offers significantly better compression than gzip for the SILVA 18-S dataset, with a compression ratio of 69.96x for panCAKES compared with 4.397 for gzip. We observe that panCAKES offers comparable compression to gzip for the set datasets, Kosarak and MovieLens 10M. For the GreenGenes 13.5 and 12.10 datasets, panCAKES offers worse compression than gzip. For the PDB-seq protein dataset, panCAKES does not offer any compression at all, as illustrated by the compression ratio of 0.61x.
Table 1: Compression Ratios.
@@ 4.3 Compression Ratios
4.4 Search
Table 2 reports the search time in seconds per query for our four algorithms on raw datasets. Table 3 reports the same figure for datasets compressed using panCAKES. In Table 4, we report the ratio of compressive search time to raw data search time for each dataset. We observe that, with the exception of the PDB-Seq dataset, panCAKES loses search speed while gaining compression. For $\rho$-NN search, the loss in search speed ranges from 1.28x to 28.26x. For repeated $\rho$-NN search, figure ranges from 1.20x to 8.99x. BFS-$k$-NN search loses speed by a factor of 0.90x to 6.50x, and DFS-$k$-NN search loses speed by a factor of 0.27x to 4.18x.
@@ 4.4 Search
Table 2: Search time in seconds per query.
@@
Table 3: Compressive Search time in seconds per query.
@@
In Table 4, we summarize the compression ratios achieved by panCAKES and the trade-off in loss of search speed.
Table 4: Summary Table of the factors by which panCAKES loses search speed while gaining compression.
@@
5 Discussion and Future Work
We have presented panCAKES, a novel approach to compression and compressive-search for big data. The approach is generic over any distance function for which the distance between two points is proportional to the memory cost of storing an encoding of one in terms of the other.
5.1 Discussion
Our results illustrate the tradeoff between compression ratios and search speed. Though panCAKES had the greatest compression ratios on the two aligned genomic datasets (SILVA and GreenGenes 12.10), these datasets exhibited relatively slow search times. Notably, panCAKES achieves  3x better compression on the SILVA dataset than gzip. While panCAKES compression is comparable to that of gzip on the two set datasets (Kosarak and MovieLens), $k$-NN search is only mildly slower on the compressed representations of these datasets than on the raw data. We stress that the primary advantage of panCAKES is that search can happen without decompressing the whole dataset, which is not possible with general purpose compression algorithms such as gzip.
@@ 5.1 Discussion
5.2 Future Work
As future work, we intend to benchmark panCAKES on additional types of data satisfying the properties outlined in Section 1. These include integer vector data under Manhattan distance, image datasets under Wasserstein distance, and radio frequency data under Dynamic Time Warping.
@@ 5.2 Future Work
Additionally, we intend to explore improvements to the compression algorithm in Section 2.2. Algorithm 2 trims the tree such that only the shallowest clusters marked for unitary compression are kept; any descendants of these clusters are discarded. This approach is greedy and may not be optimal; we may achieve better compression ratios by deferring the trimming of descendants until reaching a deeper depth in the tree.

### v2

@@ Abstract
Index Terms:
compression, search, compressive search, manifold hypothesis
I Introduction
@@
In [13], we introduced a novel hierarchical clustering algorithm, and in [4], we introduced novel algorithms for efficient, exact $k$-NN search on large datasets. Both of these approaches rely on the construction of a cluster tree, called the CLAM tree. In this paper, we extend these previous works by introducing a novel hierarchical compression algorithm, panCAKES (pan- CLAM-Accelerated K-NN Entropy Scaling Search), which uses the indexing approach from [13] and [4] to enable efficient, exact $k$-NN and $\rho$-NN search on compressed data. We stress that panCAKES’ compression is aligned with, rather than orthogonal to, speedups provided by our indexing approach. Our hierarchical cluster tree allows for both efficient compression and search on the compressed dataset. Importantly, panCAKES conducts search on compressed datasets without needing modifications to the $k$-NN and $\rho$-NN search algorithms from [4].
The motivation behind panCAKES is that many interesting datasets are too large to fit in system memory, and therefore searching them requires significant I/O. In contrast, the compressive search provided by panCAKES is designed to search datasets that do not fit in memory, while decompressing only relevant subsets of the data.
@@
II Methods
II-A Building the CLAM Tree
The compression approach presented in this paper relies on the divisive hierarchical clustering algorithm introduced in [4] and described in Algorithm 1, which constructs the CLAM tree.
@@ II-A Building the CLAM Tree
Note that Algorithm 1 does not necessarily produce a balanced tree. In fact, if we assume the manifold hypothesis–the notion that high-dimensional data collected from constrained generating phenomena typically only occupy a low-dimensional manifold within their embedding space–we expect anything but a balanced tree. The varying sampling density in different regions of the manifold and the low dimensional “shape” of the manifold itself will cause the tree to be unbalanced. The only case in which we would expect a balanced tree is if the dataset were uniformly distributed, e.g. in a $d$-dimensional hyper-sphere. We discuss additional ways that dataset structure informs properties of the cluster tree in Section IV-A.
II-B Compression
@@
Of course, for massive datasets, radii at shallow depths of the cluster tree are not small, and as discussed in Section IV-A, they are not guaranteed to decrease after the first partition. Thus, unitary compression may not be the most efficient compression approach for all clusters.
@@ III-E MovieLens-10M
IV Results
IV-A Scaling Behavior of Cluster Radii
@@
To do this, we must first introduce the concept of fractal dimensionality of a dataset. We assume the manifold hypothesis (refer to [24] or Section II-A); in other words, we assume that our dataset is embedded in a $D$-dimensional space, but that the data only occupy a $d$-dimensional manifold, where $d\ll D$. We refer to $d$ as the fractal dimensionality of the dataset.
@@
In this case, Algorithm 1 (Partition) will select a maximally distant pair of points to use as poles; without loss of generality, say we select two points at the extrema one of the $d$ axes. After one partition, the maximum pairwise distance along that axis will have been halved, i.e., bounded above by $R$. The next recursive partition will select another of the $d$ axes, and once again, the distance along that axis will be bounded above by $R$.
Thus, after at most $d$ partitions, the maximum pairwise distance along each axis will be bounded above by $R$. The overall (i.e., not restricted to one axis) maximum pairwise distance will be bounded above by $R\sqrt{2}$ for two instances that lie at the extrema of different axes. See Figure 3 for an example.
Figure 3: Scaling behavior of radii on a two dimensional disk of uniformly distributed points representing the worst-case scenario for a two-dimensional distribution. The root cluster $C_{0}$ has a center $o_{0}$ and radius radius $R_{0}$. After one application of Partition (Algorithm 1), we have a child cluster $C_{1}$, with radius $R_{1}$ and center $o_{1}$. The right triangle formed by $o_{0}$, $o_{1}$, and $y_{+}$ in $C_{1}$ shows that $R_{0}<R_{1}$. Hence, the radius of a child cluster can be larger than that of its parent. However, after another application of Algorithm 1, we have consumed both orthogonal axes, as shown in $C_{2}$. Now, clearly $R_{2}<R_{0}$.
@@
IV-B Compression Upper Bound
In this section, we provide an upper bound on the cost of compression in order to illustrate how the cost changes with respect to LFD and depth in the tree. As in Section IV-A, we assume the worst-case distribution of data and that the local fractal dimension is uniform.
To illustrate the relationship between the LFD and the cost of compression, we introduce the notion of a “stride” in the tree. If $L$ is the LFD of the dataset, we can think of a stride as an $\lceil L\rceil$-depth moving window in the tree. For example, the first stride includes all the clusters at depth 1 (root only) through depth $\lceil L\rceil$. The second stride includes the clusters at depths $\lceil L\rceil$ through $2\lceil L\rceil$. Hereafter, the ceiling is implied when LFD is non-integral; we use simply $L$ to mean $\lceil L\rceil$.
@@ IV-B Compression Upper Bound
To compute $T_{R}$, we compute the recursive cost within each stride as a function of the stride index $i$, and then sum up over the number of strides $S$. Each edge (to a child cluster from its parent) in stride $i$ represents one encoding between a pair of cluster centers. By our analysis in Section IV-A, and by our choice of a distance function under which distance values are proportional to the memory cost of the encoding, we can use $r_{i}$, the radius of the shallowest cluster in stride $i$, as an upper bound for the memory cost of each such encoding.
@@
The first factor in Equation 7 uses the radii scaling behavior guarantee from Section IV-A and the second factor is the number of non-center points in leaves at depth $S\cdot L$ (recall that we are assuming a balanced clustering).
@@
Figure 4: Compression cost breakdown assuming LFD = 3 and stride = 1. The leaf clusters within the stride (i.e., clusters at depth 4), shaded red, are compressed unitarily. Their ancestors, connected by green edges, are compressed using recursively. The recursive cost $T_{R}$ is the sum of the costs of encoding the green edges. The unitary cost $T_{U}$ is the sum of the costs of encoding all non-center points in the red clusters.
IV-C Compression Ratios
Table I reports the compression ratios of panCAKES and gzip. We observe that panCAKES offers significantly better compression than gzip for the SILVA 18-S dataset, with a compression ratio of 69.96x for panCAKES compared with 4.397 for gzip. We observe that panCAKES offers comparable compression to gzip for the set datasets, Kosarak and MovieLens 10M. For the GreenGenes 13.5 and 12.10 datasets, panCAKES offers worse compression than gzip. For the PDB-seq protein dataset, panCAKES does not offer any compression at all, as illustrated by the compression ratio of 0.61x.
TABLE I: Compression Ratios.
@@ IV-C Compression Ratios
IV-D Search
Table II reports the search time in seconds per query for our four algorithms on raw datasets. Table III reports the same figure for datasets compressed using panCAKES. In Table IV, we report the ratio of compressive search time to raw data search time for each dataset. We observe that, with the exception of the PDB-Seq dataset, panCAKES loses search speed while gaining compression. For $\rho$-NN search, the loss in search speed ranges from 1.28x to 28.26x. For repeated $\rho$-NN search, figure ranges from 1.20x to 8.99x. BFS-$k$-NN search loses speed by a factor of 0.90x to 6.50x, and DFS-$k$-NN search loses speed by a factor of 0.27x to 4.18x.
@@ IV-D Search
TABLE II: Search time in seconds per query.
@@
TABLE III: Compressive Search time in seconds per query.
@@
In Table IV, we summarize the compression ratios achieved by panCAKES and the trade-off in loss of search speed.
TABLE IV: Summary Table of the factors by which panCAKES loses search speed while gaining compression.
@@
V Discussion and Future Work
We have presented panCAKES, a novel approach to compression and compressive-search for big data. Our approach allows for efficient similarity search without decompressing the entire dataset. We achieve this result by encoding each data point as its set of differences from some representative data point; these differences can then be applied to the representative to reconstruct the original data. This reconstruction need not be applied to the entire dataset (i.e., the root of the CLAM tree), but can instead be applied to any arbitrary subtree corresponding to the result set. The approach is generic over any distance function for which the distance between two points is proportional to the memory cost of storing an encoding of one in terms of the other. In essence, panCAKES can be thought of as a simultaneous database and compression index, allowing for selective decompression based on a query.
V-A Discussion
Our results illustrate the tradeoff between compression ratios and search speed. Though panCAKES had the greatest compression ratios on the two aligned genomic datasets (SILVA and GreenGenes 12.10), these datasets exhibited relatively slow search times. Notably, panCAKES achieves  3x better compression on the SILVA dataset than gzip. While panCAKES compression is comparable to that of gzip on the two set datasets (Kosarak and MovieLens), $k$-NN search is only mildly slower on the compressed representations of these datasets than on the raw data. We stress that the primary advantage of panCAKES is that search can happen without decompressing the whole dataset, which is not possible with general purpose compression algorithms such as gzip. This essential when the dataset is vastly larger than what can fit in system memory.
@@ V-A Discussion
V-B Future Work
As future work, we intend to benchmark panCAKES on additional types of data satisfying the properties outlined in Section I. These include integer vector data under Manhattan distance, image datasets under Wasserstein distance, and radio frequency data under Dynamic Time Warping.
@@ V-B Future Work
Additionally, we intend to explore improvements to the compression algorithm in Section II-B. Algorithm 2 trims the tree such that only the shallowest clusters marked for unitary compression are kept; any descendants of these clusters are discarded. This approach is greedy and may not be optimal; we may achieve better compression ratios by deferring the trimming of descendants until reaching a deeper depth in the tree.